# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v1

@@ Abstract
When reading, we often have specific information that interests us in a text. For example, you might be reading this paper because you are curious about LLMs for eye movements in reading, the experimental design, or perhaps you only care about the question “but does it work?”. More broadly, in daily life, people approach texts with any number of text-specific goals that guide their reading behavior. In this work, we ask, for the first time, whether open-ended reading goals can be automatically decoded from eye movements in reading. To address this question, we introduce goal classification and goal reconstruction tasks and evaluation frameworks, and use large-scale eye tracking for reading data in English with hundreds of text-specific information seeking tasks. We develop and compare several discriminative and generative multimodal LLMs that combine eye movements and text for goal classification and goal reconstruction. Our experiments show considerable success on both tasks, suggesting that LLMs can extract valuable information about the readers’ text-specific goals from eye movements. [^1] Code is available anonymously here.
Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading
Cfir Avraham Hadar, Omer Shubithanks: Equal contribution., Yoav Meiri, Yevgeni Berzak Faculty of Data and Decision Sciences, Technion - Israel Institute of Technology, Haifa, Israel {kfir-hadar,shubi,meiri.yoav}@campus.technion.ac.il berzak@technion.ac.il
@@ 1 Introduction
Eye movements in reading are a key methodology for studying the cognitive basis of reading and language processing. When we read, our eyes move over the text in a saccadic manner, with prolonged periods of time (typically about 200-250ms) in which the gaze is stable, called fixations, and fast (about 30ms) transitions between fixations called saccades (Schotter and Dillon, 2025). This trajectory and the durations of the individual fixations are rich behavioral traces of the ways in which readers interact with texts and process language in real time (Rayner, 1998; Just and Carpenter, 1980). Understanding the nature and strength of the relation between eye movements, the text, and online language processing, has been a major research avenue in the psychology of reading and psycholinguistics in the past few decades, and in recent years has also been gaining increasing interest in NLP and AI (Barrett and Hollenstein, 2020; Beinborn and Hollenstein, 2024).
@@
In this work, we address a reading scenario that is prevalent in daily life but remains understudied – seeking specific information in a text. This scenario deviates from a common implicit assumption in psycholinguistics, according to which the comprehender’s goal is constant across communicative contexts: a general understanding of the linguistic input. However, in real life, readers approach texts with a variety of goals and information seeking needs. Each such goal can have a profound impact on the cognitive processes of language comprehension and the corresponding behavioral traces of eye movements over text (Radach and Kennedy, 2004; Hahn and Keller, 2023; Shubi and Berzak, 2023).
@@
Our study focuses on the following question: can arbitrary, text-specific information goals be automatically decoded from eye movements in reading? Specifically, given a single participant reading a single passage with a question in mind that they would like to answer via the text, can this question be decoded from their eye movements over the passage? This task has not been addressed in prior research and has both scientific and practical importance. Scientifically, it allows obtaining insights into the strength and nature of the relation between task conditioning and reading behavior, and more broadly, the extent to which eye movements contain information on the cognitive state of the reader. Practically, it opens the door for applications in education, content personalization, and text accessibility, that will rely on detecting information seeking behavior, its specific purpose, and the degree of its effectiveness.
In this work, we introduce two formulations of the goal decoding task, classification and reconstruction. In the classification task, given several candidate questions over a paragraph, we examine whether it is possible to predict which of these questions was provided to the reader from their eye movements over the paragraph. We examine two variants of this task with increasing difficulty: distinguishing between (i) questions about different portions of the passage, and (ii) questions over the same passage portion. In the reconstruction task, we go even further by attempting to generate the precise question. We tackle these tasks using existing and new LLM-based models that combine eye movements and text. We find that the best performing model on the classification task is able to predict the correct question with a considerable degree of success in both task variants. Our results further suggest that although the reconstruction task is extremely challenging, meaningful progress can be made even in this open-ended regime.
@@
  1. 1.
    Task We present a new cognitive state decoding challenge from eye movements in reading: decoding arbitrary information seeking goals over the content of specific texts. We instantiate this challenge as a classification task given several possible goals, and as an open-ended goal reconstruction task.
  2. 2.
    Models We adjust two discriminative models to the classification task and introduce two new generative models that combine text with eye movements in reading and can be used for both classification and reconstruction.
  3. 3.
    Evaluations We perform systematic evaluations of classification accuracy against informative baseline models in two task difficulty levels, and use several generation quality measures for the reconstruction task. For both tasks, we examine model generalization across new textual items, new readers, and the combination of both.
@@ 2 Eye Tracking Data
Our work is enabled by OneStop Eye Movements (Berzak et al., 2025, henceforth OneStop) a dataset of eye movements in reading, with 360 adult native English speakers reading Guardian articles from the 30-article OneStopQA dataset (Berzak et al., 2020). The data was collected with an EyeLink 1000 Plus eye tracker (SR Research). In each paragraph reading trial, a participant reads a paragraph on a screen, and on the following screen answers a multiple-choice reading comprehension question with 4 possible answers, without the ability to return to the paragraph. 180 participants read in an information seeking condition, where they receive the question (but not the answers) before reading the paragraph, and therefore have a specific reading goal for the upcoming paragraph. The remaining 180 participants do not receive the question ahead of time, and thus have to be ready to answer any subsequent question. In both conditions, a single participant reads one of three 10-article batches, each with 54 paragraphs. Due to the nature of the task, here, we focus on the information seeking part of the data.
In each trial, a participant receives one of three possible questions over the paragraph. Each such question has a manually annotated critical span in the paragraph which contains the information that is essential for answering the question correctly. Two of the questions share the same critical span, while the third question has a distinct critical span. This design supports a question classification task with two tiers of difficulty, distinguishing between questions with different critical spans and the more challenging variant of distinguishing between questions that share the same critical span. The critical spans are counterbalanced across the dataset such that spans that share two questions appear an equal number of times before and after spans that have one question. Figure 1 presents an example of a paragraph, its three questions, and their corresponding critical span annotations.
The information seeking portion of the data has 486 questions and 20 participants answering each question. The average paragraph length is 109 words, out of which the critical span is 32 words. The average question length is 10 words. Overall, there are 1,055,429 word tokens in the paragraphs over which eye tracking data was collected.
Figure 1: Example of a textual item from OneStopQA: a paragraph with three questions. Two of the questions have the same critical span: the paragraph segment essential for answering the question.
@@ 3 Problem Formulation
Each paragraph $P$ has two distinct critical spans, $c_{1}$ and $c_{2}$, and a set of three questions that can be answered based on the paragraph, $Q^{P}=\left\{q^{P}_{1,c_{1}},q^{P}_{2,c_{2}},q^{P}_{3,c_{2}}\right\}$. The critical information for answering $q^{P}_{1,c_{1}}$ is located in span $c_{1}$, while for both $q^{P}_{2,c_{2}}$ and $q^{P}_{3,c_{2}}$ it is in span $c_{2}$. In each experiment trial, a participant $S$ is presented with one question, $q_{S}^{P}\in Q^{P}$, before reading the paragraph $P$. The recording of the participant’s eye movements while reading the paragraph is denoted by $E^{P}_{S}$.
Question Classification
The classification task is to select which question, from a given set of questions specific to the paragraph, the participant received. We train a classifier $h$ that predicts the question that was presented to the participant prior to reading the paragraph from their eye movement recording over the paragraph:
$h(P,Q^{P},E^{P}_{S})\longrightarrow\hat{q}_{S}^{P}\in Q^{P}$
Question Reconstruction
In the reconstruction task, models are trained to generate the question that the participant received from their eye movement recording over the paragraph:
$g(P,E^{P}_{S})\longrightarrow\hat{q}_{S}^{P}$
where $g$ is a generative model whose output is a natural language question.
@@ 4 Models
We experiment with discriminative and generative modeling approaches. Figure 2 presents a schematic view of these approaches, which are further outlined below.
Figure 2: An overview of the two types of models presented in the paper. On the left, discriminative models (Haller RNN and RoBERTEye-Fixations), that score each candidate question and are evaluated on classification accuracy. On the right, generative models (DalEye-LLaVA and DalEye-Llama), that are trained to reconstruct the question presented to the reader, and are evaluated both on classification accuracy and reconstruction quality. Further details and diagrams of the models are presented in Appendix E.
@@ 4.1 Discriminative Models
Our discriminative models are adaptations of state-of-the-art encoder models for eye movements in reading to our task. Building on the approach used in Radford et al. (2019) for encoding multiple-choice reading comprehension items, we use several copies of each paragraph, where instead of combining each copy with a candidate answer, we combine it with a candidate question, and further provide the eye movements over the paragraph $\langle P,q^{P},E^{P}_{S}\rangle$. The models assign a probability to each triplet, and classification is performed by selecting the highest probability question.
Haller RNN
(Haller et al., 2022) An RNN-based model that orders words based on the fixation sequence and encodes each fixation using eye movement features. We adapt this model by appending question embeddings to the input and processing them through the RNN, followed by a classifier scoring each candidate question. The resulting architecture is depicted in Section E.3.
RoBERTEye-Fixations
(Shubi et al., 2024b) A transformer-based model which integrates fixation-level eye movement features into RoBERTa. We adapt it to incorporate candidate question embeddings and modify the classification layer to score questions. Section E.4 presents further details and a diagram of the adapted architecture.
@@ 4.2 Generative Models
We introduce two LLM-based approaches for generating the question based on the eye movements over the paragraph. Both models are trained using an autoregressive next-token prediction objective for the correct question with a cross-entropy loss. The language models’ parameters remain frozen during training, and we fine-tune only an additional LoRA (Hu et al., 2022) adapter component. As described in Section 5, we evaluate these models both on the question classification and the question reconstruction tasks.
DalEye-LLaVA
@@
is based on LLaVA-OneVision (Li et al., 2024a), a vision-language model that uses the Qwen 2-0.5B language model (Yang et al., 2024) as a backbone. The model input consists of a task prompt, the paragraph text, and eye movement feature embeddings for the paragraph words. We replace the vision encoder of LLaVA-OneVision with a novel trainable eye movement encoder, which encodes eye movement features using fully connected and convolutional layers. The eye movements embeddings are positionally aligned with the paragraph words. A diagram of the model and further details are provided in Section E.5.
DalEye-Llama
@@
In this approach, the eye movement trajectory over the paragraph is provided to the Llama 3.1 (Grattafiori et al., 2024) LLM using a textual representation. Each fixation is represented as a triplet containing the index of the fixated word, the word itself, and the fixation duration in milliseconds. This encoding preserves both the spatial locations and temporal order of the fixations within the paragraph. The resulting textual representation is combined with a prompt instructing the model to generate the question that was presented to the reader. An example of the model input for an entire fixation sequence over a paragraph, including the task prompt, is provided in Section E.6.
4.3 Baseline Models
We introduce two baseline models which use eye movements for the classification task. These models build on the following observations from prior literature: (i) readers spend more time on question-relevant than on question irrelevant portions of the text (Hahn and Keller, 2023; Malmaud et al., 2020; Shubi and Berzak, 2023), and (ii) question-relevant portions of the text have higher semantic similarity to the question compared to question-irrelevant portions of the text (Mitra et al., 2016).
Question Similarity to RT-Weighted Passage
For a participant reading a paragraph of $N$ words, we compute a paragraph embedding $\text{emb}(p)$ as a weighted average of the paragraph word embeddings, where the weights are reading times (RT) on each word:
$\text{emb}(p)=\sum_{i=1}^{N}\text{RT}(w_{i})\,\text{emb}(w_{i})$
where $\text{emb}(w_{i})$ is the contextual word embedding of word $w_{i}$, extracted from RoBERTa (Liu et al., 2019), and $\text{RT}(w_{i})$ is the Total Fixation Duration (i.e. sum of all the fixation durations) on the word normalized by total reading time of the paragraph. For each candidate question $q_{j}$, we use the embedding of the [CLS] token as $\text{emb}(q_{j})$, and compute its cosine similarity with $\text{emb}(p)$. We then predict the question with the highest similarity: $\arg\max_{j}\cos(\text{emb}(p),\text{emb}(q_{j}))$.
RT Similarity to Question-Word Similarities
Here we provide an alternative to the model above, where we compare the vector of word reading times across the paragraph to the vector of word–question similarities. Specifically, for each candidate question $q_{j}$, we first compute:
$\text{sim}(w_{i})_{q_{j}p}=\cos\bigl{(}\text{emb}(q_{j}),\text{emb}(w_{i})\bigr{)}$
where as before, $\text{emb}(w_{i})$ is a contextual word embedding from RoBERTa, and $\text{emb}(q_{j})$ is the question [CLS] embedding, yielding a length-$N$ vector $\text{sim}_{q_{j}p}$. We select the question maximizing the cosine similarity between this vector and $\text{RT}(p)$, the length-$N$ vector of Total Fixation Durations for the paragraph words: $\arg\max_{j}\cos\bigl{(}\text{sim}_{q_{j}p},\,\text{RT}(p)\bigr{)}$.

### v2

@@ Abstract
When reading, we often have specific information that interests us in a text. For example, you might be reading this paper because you are curious about LLMs for eye movements in reading, the experimental design, or perhaps you wonder “This sounds like science fiction. Does it actually work?”. More broadly, in daily life, people approach texts with any number of text-specific goals that guide their reading behavior. In this work, we ask, for the first time, whether open-ended reading goals can be automatically decoded solely from eye movements in reading. To address this question, we introduce goal decoding tasks and evaluation frameworks using large-scale eye tracking for reading data in English with hundreds of text-specific information seeking tasks. We develop and compare several discriminative and generative multimodal text and eye movements LLMs for these tasks. Our experiments show considerable success on the task of selecting the correct goal among several options, and even progress towards free-form textual reconstruction of the precise goal formulation. These results open the door for further scientific investigation of goal driven reading, as well as the development of educational and assistive technologies that will rely on real-time decoding of reader goals from their eye movements. [^1] Code is available anonymously here.
@@ 1 Introduction
Eye movements in reading are a key methodology for studying the cognitive basis of reading and language processing. When we read, our eyes move over the text in a saccadic manner, with prolonged periods of time in which the gaze is stable, called fixations, and fast transitions between fixations called saccades (Schotter & Dillon, 2025). This trajectory contains rich behavioral traces of the ways in which readers interact with texts and process language in real time (Rayner, 1998; Just & Carpenter, 1980). Understanding the nature and strength of the relation between eye movements, the text, and online language processing has been a major research avenue in the psychology of reading in the past few decades, and in recent years has also been gaining increasing interest in NLP and ML (Barrett & Hollenstein, 2020; Reich et al., 2025).
@@
In this work, we investigate a reading scenario that is prevalent in daily life but remains understudied – seeking specific information in a text. This scenario deviates from a common implicit assumption in psycholinguistics, according to which the comprehender’s goal is constant across communicative contexts: a general understanding of the linguistic input. In practice, readers approach texts with a variety of goals and information seeking needs. Each such goal can have a profound impact on the cognitive processes of language comprehension and the corresponding behavioral traces of eye movements over the text (Radach & Kennedy, 2004; Kaakinen & Hyönä, 2010; Hahn & Keller, 2023; Shubi & Berzak, 2023).
@@
Our study focuses on the following question: can arbitrary, text-specific information goals be automatically decoded from eye movements in reading? Specifically, given a single participant reading a single passage with a question in mind that they would like to answer via the text, can this question be decoded from their eye movements over the passage? This task has not been addressed in prior research and has both scientific and practical importance. Scientifically, it allows obtaining insights into the strength and nature of the relation between task conditioning and reading behavior, and more broadly, the extent to which eye movements contain information on the cognitive state of the reader. Practically, it opens the door for applications in education, content personalization, and text accessibility, which will rely on detecting information seeking behavior, its specific purpose, and the extent to which this behavior is effective.
@@
Tasks: We present a new cognitive state decoding challenge from eye movements in reading: decoding arbitrary information seeking goals over the content of specific texts. We instantiate this challenge as a goal selection task given several possible goals, and as an even more challenging open-ended goal reconstruction task where the information-seeking goal has to be generated.
Models: We introduce two strong baseline models and adjust two state-of-the-art discriminative models that combine text with eye movements for the goal selection task. We further develop two new generative text and eye movements multimodal LLMs that can be used for both goal selection and goal reconstruction.
Evaluations: We perform systematic evaluations of selection accuracy against informative baseline models at two levels of task difficulty and utilize several generation quality measures for the reconstruction task. For both tasks, we examine model generalization across new texts, new readers, and the combination of both.
@@ 2 Eye Tracking Data
We use OneStop (Berzak et al., 2025), a public eye tracking dataset collected with a high-end EyeLink 1000 Plus eye tracker. It comprises 360 adult native English speakers reading Guardian articles from the OneStopQA multiple choice reading comprehension dataset (Berzak et al., 2020; Vajjala & Lučić, 2018). In each experimental trial, a participant reads a paragraph on a screen and, on the following screen, answers a multiple-choice reading comprehension question, without the ability to return to the paragraph. 180 participants read in an information seeking condition, where they receive the question (but not the answers) before reading the paragraph, and therefore have a specific reading goal for the upcoming paragraph. The remaining 180 participants do not receive the question ahead of time, and thus have to be ready to answer any subsequent question. Here, we use the information seeking part of OneStop, which is uniquely suited for our study.
Figure 1: Example of a text from OneStop, consisting of a paragraph and three questions. Two of the questions have the same critical span: the paragraph segment essential for answering the question.
For each paragraph, a participant receives one of three possible questions over the paragraph. Each such question has a manually annotated critical span in the paragraph, which contains the information that is essential for answering the question correctly. Two of the questions share the same critical span, while the third question has a distinct critical span. This design supports a question selection task with two tiers of difficulty, distinguishing between questions with different critical spans and the more challenging variant of distinguishing between questions that share the same critical span. Figure˜1 presents an example of a paragraph, its three questions, and their corresponding critical span annotations.
The data has 486 questions over 162 paragraphs, with 20 participants answering each question. The mean question length is 10 words. The mean paragraph length is 109 words (5.8 sentences), of which the critical span is 32 words. Overall, there are 1,055,429 paragraph word tokens for which eye tracking data was collected in the information seeking regime.
@@ 3 Problem Formulation
The general problem we address is using eye movements over a text (in our setting, a text is always a paragraph) to predict which question a participant received prior to reading that text. Each text $T$ has three text specific questions, $Q_{T}=\left\{q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}\right\}$. The critical information for answering $q_{1,c_{1}}$ is located in the text in span $c_{1}$, while for both $q_{2,c_{2}}$ and $q_{3,c_{2}}$ it is in span $c_{2}$. In each experiment trial, a participant $P$ is presented with one question, $q_{P}\in Q_{T}$, before reading the text. The recording of the participant’s eye movements while reading the text is denoted by $E_{P,T}$.
We propose two task variants as follows.
Question Selection In the selection task, a classifier $h$ is trained to select which question from $Q_{T}$ the participant received, from their eye movement recording over the text:
$h(T,Q_{T},E_{P,T})\longrightarrow\hat{q}_{P}\in Q_{T}$
Question Reconstruction In the free-form reconstruction task, a generative model $g$ is trained to generate the question that the participant received from their eye movement recording over the text:
$g(T,E_{P,T})\longrightarrow\hat{q}_{P}$
@@ 4 Models
We introduce discriminative and generative models, presented schematically in Figure˜2.
(a) Discriminative models (Section˜4.1)
(b) Generative models (Section˜4.2)
Figure 2: Two model types for decoding the question presented to a participant before reading the text, from their eye movements while reading the text. (a) Discriminative models that score candidate questions and are evaluated on question selection accuracy. (b) Generative models trained to reconstruct the question presented to the reader, and evaluated on question selection accuracy (via log-likelihood assignment to each candidate question) and free-form question reconstruction quality.
@@ 4.1 Discriminative Models
Reading-Time Informed Embedding Similarity Models
We develop two simple baseline models which build on the following observations from prior literature: (i) readers tend to spend more time on question-relevant than on question-irrelevant portions of the text (Hahn & Keller, 2023; Malmaud et al., 2020; Shubi & Berzak, 2023), and (ii) question-relevant text segments tend to have higher semantic similarity to the question compared to question-irrelevant segments (Mitra et al., 2016).
(a) Question Similarity to RT-Weighted Passage
(b) RT Similarity to Question-Word Similarities
Figure 3: Reading-Time Informed Embedding Similarity models for the question selection task.
Question Similarity to RT-Weighted Passage (Figure˜3(a)) For a participant reading a text of $N$ words, we compute a text embedding, $\text{emb}(T)$, as a weighted average of its contextualized word embeddings, where the weights are speed-normalized word reading times, $RT(w_{i})$. Formally: $\text{emb}(T)=\sum_{i=1}^{N}\text{RT}(w_{i})\,\text{emb}(w_{i})$. The embeddings for word $w_{i}$, $\text{emb}(w_{i})$, are extracted from RoBERTa (Liu et al., 2019), and $\text{RT}(w_{i})$ is the word’s Total Fixation Duration (i.e., sum of all the fixation durations) normalized by the text’s total reading time. For each candidate question $q_{j}$, we use its [CLS] token embedding as $\text{emb}(q_{j})$, and compute its cosine similarity with $\text{emb}(T)$. We then predict the question with the highest similarity:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos(\text{emb}(T),\text{emb}(q_{j}))$
RT Similarity to Question-Word Similarities (Figure˜3(b)) An alternative to the model above. Compares the vector of word reading times for the text to the vector of word–question similarities. Specifically, for each candidate question $q_{j}$ and text word $w_{i}$, we first compute $\text{sim}(q_{j},w_{i})=\cos\bigl(\text{emb}(q_{j}),\text{emb}(w_{i})\bigr)$, where as before, $\text{emb}(w_{i})$ is a contextual word embedding from RoBERTa, and $\text{emb}(q_{j})$ is the question [CLS] embedding, yielding a length-$N$ vector of similarities $\text{sim}(q_{j},T)=[\text{sim}(q_{j},w_{1}),\text{sim}(q_{j},w_{2}),...,\text{sim}(q_{j},w_{n})]$. We select the question maximizing the cosine similarity between this vector and $\text{RT}(T)$, the length-$N$ vector of speed-normalized Total Fixation Durations for the text words:
${\arg\max}_{j\in\{1,_{\cdots},|Q_{T}|\}}\cos\bigl(\text{sim}(q_{j},T),\,\text{RT}(T)\bigr)$
State-of-the-art Neural Models
We further adapt two state-of-the-art encoder models for eye movements in reading to our task. Building on the approach used in Radford et al. (2019) for encoding multiple-choice reading comprehension items, we use several copies of each text, where instead of combining each copy with a candidate answer, we combine it with a candidate question, and further provide the eye movements over the text $\langle T,q\in Q_{T},E_{P,T}\rangle$. The models assign a probability to each triplet and select the highest probability question.
Haller RNN (Haller et al., 2022) An RNN-based model that orders word embeddings based on the fixation sequence and encodes each fixation using eye movement features. We adapt this model by appending question embeddings to the input and processing them through the RNN, followed by a classifier scoring each candidate question. Additional details on the model, including a diagram of the modified architecture and the used eye movement features, are presented in Section˜A.1.
RoBERTEye-Fixations (Shubi et al., 2024) A transformer-based model which integrates fixation-level eye movement features into RoBERTa (Liu et al., 2019). An eye movements feature vector for each fixation is first projected to the dimension of the word embeddings. These fixation vectors are then aligned with their corresponding words via the addition of position embeddings and finally concatenated with the word sequence. We adapt this model to incorporate candidate question embeddings and modify the classification layer to score questions. Section˜A.2 provides further details, including a model diagram and the eye movement features that the model uses.
@@ 4.2 Generative Models
We introduce two LLM-based approaches for generating the question based on the eye movements over the text. Both models are fine-tuned using an autoregressive next-token prediction objective for the correct question with a cross-entropy loss. During fine-tuning, the language models’ parameters remain frozen, and only an additional LoRA (Hu et al., 2022) adapter component is trained. As described in Section˜5, we evaluate these models both on the question selection and the question reconstruction tasks.
@@
DalEye-LLaVA is based on LLaVA-OneVision (Li et al., 2024), a vision-language model that uses the Qwen 2-0.5B language model (Yang et al., 2024) as a backbone. The model input consists of a task prompt, the text, and eye movement feature embeddings for the text words. We replace the vision encoder of LLaVA-OneVision with a novel trainable eye movement encoder, which encodes eye movement features using fully connected and convolutional layers. The eye movements embeddings are positionally aligned with the text words. Additional details, including a model diagram and eye movement features, are provided in Section˜B.1.
@@
DalEye-Llama In this approach, we provide the task and the eye movement trajectory of a participant over the text to the Llama 3.1 language model (Grattafiori et al., 2024) in textual form. The model is finetuned to reconstruct the true question given this input. We use a fixation-based eye movements representation, where each fixation is represented as a tuple containing the index of the fixated word, the word itself, the fixation duration in milliseconds, and the direction of the next saccade. This encoding preserves both the spatial locations and temporal order of the fixations on the text. It follows a prompt describing the task and instructing the model to generate the question that was presented to the reader. In Section˜B.2, we provide an example of the model input for a complete fixation sequence over a paragraph. We further report experimental results for a differently worded task prompt, as well as two alternative representations of the eye movement sequence, one based on word-level features and another combining word- and fixation-level features.